A race car makes one lap around a track of radius 60 m in 6 s. What is the average velocity of the car as it goes around the circular track? Round your answer to one decimal place.

1
Expert's answer
2022-02-17T04:36:12-0500

Given:

"r=60\:\rm m"

"T=6\:\rm s"


The average velocity of the car

"v=\frac{l}{t}=\frac{2\pi r}{T}"

"v=\frac{2\pi *60}{6}=20\pi\:\rm m/s=62.8\: m/s"